OpenStudy (anonymous):

2x^3-12x^2+ 18x-108/x-6

OpenStudy (anonymous):

A. 4x2 + 9x B. 2x2 + 18 C. 3x + 5 D. 3x + 2

OpenStudy (anonymous):

$$\frac{2x^3-12x^2+18x-108}{x-6}=\frac{2x^2(x-6)+18(x-6)}{x-6}=\frac{(2x^2+18)(x-6)}{x-6}$$

OpenStudy (anonymous):

it is b?